OpenStudy (anonymous):

Solve this Proportion: 5t-3/-2=t+3/2

OpenStudy (anonymous):

given 5t-(3/-2)=t+3/2 or, 4t+3/2=3/2 or, t=0

OpenStudy (anonymous):

Can you write the equation more clearly? Cause it can be \[(5t-3)/-2=(t+3)/2\] or others too...

OpenStudy (anonymous):

The way I wrote it and the way you wrote it can be answered the same.

OpenStudy (anonymous):

If the three was negative, I would help you out by putting (-3) :P Just saying

OpenStudy (anonymous):

Then wouldn't you write it as 5x+(3/2)? ==

OpenStudy (anonymous):

5t - 3 / -2 = t + 3 / 2 t=?

OpenStudy (anonymous):

\[(5t-3)/-2=(t+3)/2\]\[3-5t=t+3\]\[6t=0\]\[t=0\]

OpenStudy (anonymous):

The first step I did was simplification.